# A new generation of effective core potentials: Selected lanthanides and heavy elements II

### v1

@@ Abstract
We present a new set of correlation-consistent effective core potentials (ccECPs) for selected heavy , , , and -block elements significant in materials science and chemistry (Rb, Sr, Cs, Ba, In, Sb, Pb, Ru, Cd, La, Ce, and Eu). The ccECPs are designed using minimal Gaussian parameterization to achieve smooth and bounded potentials. They are expressed as a combination of averaged relativistic effective potentials (AREP) and effective spin-orbit (SO) terms, developed within a relativistic coupled-cluster framework. The optimization is driven by correlated all-electron (AE) atomic spectra, norm-conservation, and spin-orbit splittings, with considerations for plane wave cut-offs to ensure accuracy and viability across various electronic configurations. Transferability of these ccECPs is validated through testing on molecular oxides and hydrides, emphasizing discrepancies in molecular binding energies across a spectrum of bond lengths and electronic environments. The ccECPs demonstrate excellent agreement with AE reference calculations, attaining chemical accuracy in bond dissociation energies and equilibrium bond lengths, even in systems characterized by substantial relativistic and correlation effects. These ccECPs provide accurate and transferable framework for valence-only calculations.
@@ I Introduction
Effective core potentials (ECPs) and closely related pseudopotentials are essential tools in electronic structure calculations, enabling us to focus on valence electrons by eliminating core electrons and corresponding core energy scales. This partitioning assumes that cores have only marginal influence on the properties of valence electronic states that determine behavior of chemical and condensed matter systems such as bonding, reactivity, magnetic effects and many others. In an ECP construction the core states are substituted by potentials that mimic the original effects of the core on valence electrons, significantly simplifying the subsequent, valence-only calculations. This is particularly important for heavy elements where core states modifications due to relativistic effects complicate the electron-electron correlation effects making thus all-electron (AE) calculations more challenging. In such scenarios, ECPs become almost indispensable as they improve efficiency while effectively capture major parts of core-valence and core-core correlations that could otherwise be neglected, for example, in frozen core approaches.
@@
Historically, ECP constructions have been built around self-consistent approaches such as Hartree-Fock/Dirac-Fock (HF/DF) and variety of Density Functional Approximations (DFAs) by reproducing one-particle atomic eigenvalues, norm/charge conservation outside the cores, and total energy differences between ground and excited states.
Notably, over the past few years, we have introduced correlation-consistent ECPs (ccECPs) with the major effort in boosting their overall fidelity to AE Hamiltonians by employing correlated, many-body approaches simultaneously for both AE and ECP construction procedures. One of our construction targets is reproducing many-body atomic energy differences (excitations and ionizations) for valence part of the atomic spectra. Another aim is transferability that has been significantly enhanced by reproducing the binding curves of hydride and oxide molecules which probe covalent and ionic bonds as they compare with AE, fully correlated equivalents. In addition, we have tried to reach the limits of simple semi-local gaussian-based forms that provide convenient use in many codes.
@@
Our previous careful testing revealed and quantified also systematic errors for cases with large shallow cores and resulting small number of valence electrons for example, for Na[Ne] and Mg[Ne] (here the element in the bracket implies the size of the ionic core). Occasionally, we have found significant discrepancies with regard to AE references and for these cases we tested smaller cores such as Na[He] which showed excellent fidelity with AE results. Further developments have included Kleinman-Bylander transformation for use in plane wave codes, softening the ECPs for lower plane wave cut-offs in codes with periodic boundary conditions, and providing the basis sets appropriate for correlated calculations. A more extensive and detailed descriptions can be found in Refs.[1, 2, 3, 4, 5, 6, 7].
In this continuing work, we expand ccECP library to include the and metals (Rb, Sr, Cs, Ba), and post-transition metals and metalloids (In, Sb, Pb), as well as technologically important transition metals (Ru, Cd) and lanthanides (La, Ce, Eu). The construction methodology emphasizes the use of a minimal parameterization with a simple Gaussian functional form, ensuring smooth and bounded potentials. We also prioritize the transferability of these ECPs by validating their performance across multiple bonding environments as done previously by testing oxide and hydride binding curves for a range of bond lengths.
@@ II.2 Optimization procedure
The optimization process begins with constructing the AREP portion, as this component is essential for achieving the required accuracy and transferability of the ccECP.
Once this foundation is established, we proceed to optimize the SO terms, ensuring that the full relativistic effects are accurately incorporated into the model.
@@
We targeted the same bonding configuration prescribed by Dolg and company in their work on the lanthanides[30].
@@ III.1 Selected elements
In this section, we present results for rubidium (Rb) and strontium (Sr), both using a 28-electron core [[Ar]4] with , , and non-local channels, consistent with our previous ccECP work[2]. For Sr, we developed two ccECP variants: small-core [[Ar]4] and large-core [Kr].
| Atom | Hamiltonian | Atom | Hamiltonian | |||||||||||
| Rb | 9 | AREP | 0 | 2 | 5.462652 | 89.500251 | Sr | 10 | AREP | 0 | 2 | 7.439649 | 131.279952 | |
| 0 | 2 | 2.056983 | 0.319839 | 0 | 2 | 3.447122 | 19.222463 | |||||||
| 1 | 2 | 4.588244 | 48.545978 | 1 | 2 | 4.744598 | 77.982723 | |||||||
| 1 | 2 | 1.032036 | 0.790029 | 1 | 2 | 8.237021 | 15.520051 | |||||||
| 2 | 2 | 3.413959 | 26.206712 | 2 | 2 | 3.183727 | 29.113266 | |||||||
| 2 | 2 | 1.024060 | 0.674894 | 2 | 2 | 7.748102 | 5.427551 | |||||||
| 3 | 1 | 3.500301 | 9.000000 | 3 | 1 | 9.273368 | 10.000000 | |||||||
| 3 | 3 | 4.112875 | 31.502713 | 3 | 3 | 9.163478 | 92.733680 | |||||||
| 3 | 2 | 3.366936 | -10.463214 | 3 | 2 | 10.340438 | -4.567422 | |||||||
| 3 | 2 | 1.707218 | -2.167673 | 3 | 2 | 9.152969 | -95.475410 | |||||||
| SO | 1 | 2 | 4.388251 | -39.040315 | SO | 1 | 2 | 7.222917 | -58.876406 | |||||
| 1 | 2 | 4.327577 | 39.039369 | 1 | 2 | 7.171633 | 58.880316 | |||||||
| 1 | 2 | 1.357223 | -0.331685 | 1 | 2 | 3.003530 | -9.875334 | |||||||
| 1 | 2 | 1.230709 | 0.437134 | 1 | 2 | 2.833933 | 9.735187 | |||||||
| 2 | 2 | 3.409705 | -10.483338 | 2 | 2 | 6.319959 | -11.907111 | |||||||
| 2 | 2 | 3.418975 | 10.482131 | 2 | 2 | 6.391156 | 11.906538 | |||||||
| 2 | 2 | 1.022484 | -0.279419 | 2 | 2 | 1.767711 | -2.201784 | |||||||
| 2 | 2 | 1.027724 | 0.260697 | 2 | 2 | 1.637060 | 1.928907 | |||||||
| Sr | 2 | AREP | 0 | 2 | 0.825002 | 15.386176 | ||||||||
| 1 | 2 | 0.397788 | 5.077782 | |||||||||||
| 2 | 2 | 0.453259 | -2.256103 | |||||||||||
| 3 | 1 | 3.000056 | 2.000000 | |||||||||||
| 3 | 3 | 3.002302 | 6.000111 | |||||||||||
@@
Rb has closed and shells and a half-filled shell. This configuration with 9 valence electrons lends itself to developing a robustly performing ECP. From Figs. 1 and 4 we see the legacy ECPs and our ccECP perform well within chemical accuracy in the atomic and molecular spectra. Only, LANL2 deviates significantly from the AE atomic spectrum and this discrepancy is amplified in the molecular properties. Therefore, in order to provide a better comparison between other ECPs, LANL2 has been omitted from the molecular binding curves as it was out of bounds. For the [[Ar]3] core, there are few ECPs available to compare against and our ccECP shows a slight overall improvement specifically close to the equilibrium bond lengths and in the LMAD criterion.
@@
For Sr, the small-core [[Ar]3] ccECP performs well in the atomic spectrum, matching the accuracy of CRENBL and UC. For SrH hydride (Fig. 5), which has a relatively simple electronic structure, the ccECP achieves chemical accuracy comparable to that of legacy ECPs. For SrO, which is a more challenging system, the ccECP shows minimal discrepancies across most bond lengths and matches the performance of UC.
Compared to the small-core, the large-core [Kr] variant shows noticeable differences in isospectrality (Figs. 1-3) and in molecular binding curves (Fig. 6). In the atomic spectrum, the ccECP outperforms all legacy ECPs and the UC results, with MAD values below 0.1 eV. In the molecules, despite the challenge of having only two valence electrons, our ccECP is the only ECP to perform consistently within the bounds of chemical accuracy across many of the geometries tested. The ccECP offers negligible discrepancies at the equilibrium bond length, approaching the limits of this construction. Although the large core variant is efficient, the small core [[Ar]4] remains a viable for studies that require proper description of semi-core levels.
@@ III.2 Selected elements
The simple electronic structure of Cs bonding is reflected in CsH and CsO molecules in Fig. 7 where most of the ECPs are well within the chemical accuracy regime. The ccECP shows noticeable improvements at the equilibrium bond length (Fig. 7). Similar to the case of Rb, LANL2 has been omitted from the molecular binding curves due its large discrepancies. Furthermore, for the ccECP we observe a systematic improvement in the atomic spectrum compared to the legacy ECPs, as shown in Figs. 1-3
@@
Despite Ba having no occupied electrons in its ground state, the channel is crucial for capturing subtle electronic interactions and relativistic effects, ensuring a more precise representation of this atom in complex systems. During optimization, we sought a balance between accurately replicating the AE atomic spectrum and maintaining robust performance across various chemical environments. Our primary focus was to prioritize precision in the atomic spectrum, accepting a minor trade-off in molecular binding curves (Fig. 8(b)) for BaO at compressed bond lengths. The ccECP demonstrates consistent results within chemical accuracy across a wide array of bond lengths, particularly at the equilibrium geometry. We have successfully mitigated the overbinding issue at compressed bond lengths, achieving higher accuracy compared to several legacy ECPs and notably outperforming even the UC method in this critical region. This approach, in our view, represents the optimal balance, maximizing the ccECP’s overall efficacy and reliability.
@@ III.3 Selected and elements
For antimony (Sb) there are nominally 5 valence electrons, which makes it rather difficult to capture accurate properties for all the tested possibilities. The results presented in Figs. 1-3 indicate that our ccECP performs on par or better than most of the legacy ECPs and the UC method. While it is possible to further improve the MADs through optimization, it comes at the cost of negatively impacting the accuracy of the binding curve in SbO. The optimal compromise is encouraging in comparison with the legacy constructions, see Fig. 11, while also providing lower MADs and LMADs. This required significant interventions into the optimization process of ccECP parameters, in particular, special care given to the and local channels. This resulted in the discrepancies being well below chemical accuracy near equilibrium bond lengths, 11.
@@
For lead (Pb), converging calculations at the CCSD(T) level required the use of unrestricted CCSD(T) and adjustments of several optimization parameters in order to aid the convergence. In the atomic spectrum, ccECP shows a significant improvement over legacy ECPs and the UC method, achieving the best overall metrics and approaching chemical accuracy.Molecular calculations (Fig. 12) proved more challenging for the ECPs, particularly for PbO, where legacy ECPs and the UC method exhibit discrepancies almost an order of magnitude or more beyond the chemical accuracy. In PbH, our ccECP balances performance across bond lengths, addressing overbinding at the dissociation limit (seen in MDFSTU, MHFSTU, MWBSTU, and SBKJC) and underbinding at compressed bond lengths (observed with UC and CRENBL). For PbO, the ccECP demonstrates consistent improvement, maintaining chemical accuracy across most geometries and at the equilibrium configuration, while significantly reducing errors at compressed bond lengths.

### v2

@@ Abstract
We present a new set of correlation-consistent effective core potentials (ccECPs) for selected heavy , , , and -block elements significant in materials science and chemistry (Rb, Sr, Cs, Ba, In, Sb, Pb, Ru, Cd, La, Ce, and Eu). The ccECPs are designed using minimal Gaussian parameterization to achieve smooth and bounded potentials. They are expressed as a combination of averaged relativistic effective potentials (AREP) and effective spin-orbit (SO) terms, developed within a relativistic coupled-cluster framework. The optimization is driven by correlated all-electron (AE) atomic spectra, norm-conservation, and spin-orbit splittings, with considerations for plane wave cut-offs to ensure accuracy and viability across various electronic configurations. The transferability of these ccECPs is validated through testing on molecular oxides and hydrides, emphasizing discrepancies in molecular binding energies across a spectrum of bond lengths and electronic environments. The ccECPs demonstrate excellent agreement with AE reference calculations, attaining chemical accuracy in bond dissociation energies and equilibrium bond lengths, even in systems characterized by substantial relativistic and correlation effects. These ccECPs provide an accurate and transferable framework for valence-only calculations.
@@ I Introduction
Effective core potentials (ECPs) and closely related pseudopotentials are essential tools in electronic structure calculations, enabling us to focus on valence electrons by eliminating core electrons and the corresponding core energy scales. This partitioning assumes that cores have only marginal influence on the properties of valence electronic states that determine behavior of chemical and condensed matter systems such as bonding, reactivity, magnetic effects, and many others. In an ECP construction, the core states are substituted by potentials that mimic the original effects of the core on the valence electrons, significantly simplifying the subsequent valence-only calculations. This is particularly important for heavy elements from the fifth and sixth rows of the periodic table, where core states modifications due to relativistic effects complicate the electron-electron correlation effects making thus all-electron (AE) calculations more challenging. In such scenarios, ECPs become almost indispensable as they improve efficiency while effectively capture major parts of core-valence and core-core correlations that could otherwise be neglected, for example, in frozen core approaches.
Notably, over recent years, we have developed correlation-consistent ECPs (ccECPs), enhancing their overall fidelity to AE Hamiltonians through correlated many-body methodologies applied consistently to both AE and ECP construction [1, 2, 3, 4, 5, 6, 7]. Primary objectives include reproducing valence many-body atomic energy differences (excitations and ionizations), ensuring finite wavefunction cusp conditions at the origin, and improving transferability. The latter has been achieved by matching binding curves of hydride and oxide molecules to their fully correlated AE counterparts, thereby probing covalent and ionic bonding. We have concurrently explored the performance limits of simple semi-local Gaussian-based forms, which enable convenient implementation across computational frameworks. The reliability of ccECPs has been rigorously established through extensive validation both in our efforts and by independent research groups, who have consistently leveraged its pragmatism in diverse many-body calculations—from solids and large molecular complexes to neural-network-based quantum simulations—demonstrating robust transferability across chemical environments[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
In contrast, legacy ECPs (referring to previous-generation constructions) were typically built around self-consistent approaches such as Hartree-Fock/Dirac-Fock (HF/DF) and a variety of Density Functional Approximations, focusing on reproducing one-particle atomic eigenvalues, norm/charge conservation outside the cores, and total energy differences between ground and excited states [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 27, 28, 43, 35, 44, 45, 46, 47, 48, 49].
@@
Our previous careful testing revealed and quantified also systematic errors for cases with large shallow cores and resulting small number of valence electrons for example, for Na[Ne] and Mg[Ne] (here the element in the bracket implies the size of the ionic core). Occasionally, we have found significant discrepancies with regard to AE references and for these cases we tested smaller cores such as Na[He] which showed excellent fidelity with AE results. Further developments have included Kleinman-Bylander transformation for use in plane wave codes, softening the ECPs for lower plane wave cut-offs in codes with periodic boundary conditions, and providing the basis sets appropriate for correlated calculations.
In this continuing work, we expand ccECP library to include elements specifically requested by the user community through the pseudopotential library[50]: and metals (Rb, Sr, Cs, Ba), and post-transition metals and metalloids (In, Sb, Pb), as well as technologically important transition metals (Ru, Cd) and lanthanides (La, Ce, Eu). This community-directed expansion addresses urgent computational needs across frontier applications including heterogeneous catalysis, advanced materials design, and correlated electron systems in condensed matter physics. Maintaining our methodology’s core principles, we employ minimal parameterization with simple Gaussian functional forms to ensure smooth, bounded potentials. Transferability is further verified through rigorous validation across diverse bonding environments, including oxide and hydride binding curves that span comprehensive bond-length regimes.
@@ II.2 Optimization procedure
In order to find the optimal values for the Gaussian parameters denoted with Greek letters in Eqs.(6–8) we follow an optimization process that begins with constructing the AREP portion, as this component is essential for achieving the required accuracy and transferability of the ccECP. Once this foundation is established, we proceed to optimize the SO terms, ensuring that the full relativistic effects are accurately incorporated into the model.
@@
We targeted the same bonding configuration prescribed by Dolg and coworkers in their work on the lanthanides[43].
@@ III.1 Selected elements
In this section, we present results for rubidium (Rb) and strontium (Sr), both using a 28-electron core [[Ar]4] with , , and non-local channels, consistent with our previous ccECP work[2]. For Sr, we developed two variants of ccECP: small-core [[Ar]4] and large-core [Kr].
| Atom | Hamiltonian | Atom | Hamiltonian | ||||||||||||
| Rb | 9 | AREP | 0 | 2 | 5.462653 | 89.500251 | 89.500251 | Sr | 10 | AREP | 0 | 2 | 7.439649 | 131.279952 | |
| 0 | 2 | 2.056984 | 0.319840 | 0.319839 | 0 | 2 | 3.447122 | 19.222463 | |||||||
| 1 | 2 | 4.588244 | 48.545979 | 48.545978 | 1 | 2 | 4.744598 | 77.982723 | |||||||
| 1 | 2 | 1.032037 | 0.790029 | 0.790029 | 1 | 2 | 8.237021 | 15.520051 | |||||||
| 2 | 2 | 3.413959 | 26.206712 | 26.206712 | 2 | 2 | 3.183727 | 29.113266 | |||||||
| 2 | 2 | 1.024061 | 0.674894 | 0.674894 | 2 | 2 | 7.748102 | 5.427551 | |||||||
| 3 | 1 | 3.500301 | 9.000000 | 9.000000 | 3 | 1 | 9.273368 | 10.000000 | |||||||
| 3 | 3 | 4.112875 | 31.502713 | 31.502713 | 3 | 3 | 9.163478 | 92.733680 | |||||||
| 3 | 2 | 3.366937 | -10.463215 | -10.463214 | 3 | 2 | 10.340438 | -4.567422 | |||||||
| 3 | 2 | 1.707218 | -2.167674 | -2.167673 | 3 | 2 | 9.152969 | -95.475410 | |||||||
| SO | 1 | 2 | 4.388251 | -39.040315 | -39.040315 | SO | 1 | 2 | 7.222917 | -58.876406 | |||||
| 1 | 2 | 4.327577 | 39.039369 | 39.039369 | 1 | 2 | 7.171633 | 58.880316 | |||||||
| 1 | 2 | 1.357223 | -0.331685 | -0.331685 | 1 | 2 | 3.003530 | -9.875334 | |||||||
| 1 | 2 | 1.230710 | 0.437135 | 0.437134 | 1 | 2 | 2.833933 | 9.735187 | |||||||
| 2 | 2 | 3.409706 | -10.483339 | -10.483338 | 2 | 2 | 6.319959 | -11.907111 | |||||||
| 2 | 2 | 3.418975 | 10.482132 | 10.482131 | 2 | 2 | 6.391156 | 11.906538 | |||||||
| 2 | 2 | 1.022484 | -0.279420 | -0.279419 | 2 | 2 | 1.767711 | -2.201784 | |||||||
| 2 | 2 | 1.027724 | 0.260697 | 0.260697 | 2 | 2 | 1.637060 | 1.928907 | |||||||
| Sr | 2 | AREP | 0 | 2 | 0.825002 | 15.386176 | |||||||||
| 1 | 2 | 0.397788 | 5.077782 | ||||||||||||
| 2 | 2 | 0.453259 | -2.256103 | ||||||||||||
| 3 | 1 | 3.000056 | 2.000000 | ||||||||||||
| 3 | 3 | 3.002302 | 6.000111 | ||||||||||||
@@
Rb has closed and shells and a half-filled shell. This configuration with 9 valence electrons lends itself to developing a robustly performing ECP. From Figs. 1 and 4 we see that the legacy ECPs and our ccECP perform well within chemical accuracy in the atomic and molecular spectra. Only, LANL2 deviates significantly from the AE atomic spectrum and this discrepancy is amplified in the molecular properties. Therefore, in order to provide a better comparison between other ECPs, LANL2 has been omitted from the molecular binding curves as it was out of bounds. For the [[Ar]3] core, there are few ECPs available to compare against, and our ccECP shows a slight overall improvement specifically close to the equilibrium bond lengths and in the LMAD criterion.
@@
For Sr, the small-core [[Ar]3] ccECP performs well in the atomic spectrum, matching the accuracy of CRENBL and UC. For SrH hydride (Fig. 5), which has a relatively simple electronic structure, the ccECP achieves a chemical accuracy comparable to that of legacy ECPs. For SrO, which is a more challenging system, the ccECP shows minimal discrepancies ( 0.02 ev) across most bond lengths and matches the performance of UC.
Compared to the small-core, the large-core [Kr] variant shows noticeable differences in isospectrality (Figs. 1-3) and in molecular binding curves (Fig. 6). In the atomic spectrum, the ccECP outperforms all legacy ECPs and the UC results, with MAD values below 0.1 eV. In molecules, despite the challenge of having only two valence electrons, our ccECP is the only ECP that perform consistently within the bounds of chemical accuracy across many of the geometries tested. The ccECP offers negligible discrepancies at the equilibrium bond length, approaching the limits of this construction. Although the large core variant is efficient, the small core [[Ar]4] remains viable for studies that require a proper description of semi-core levels.
@@ III.2 Selected elements
The simple electronic structure of Cs bonding is reflected in the CsH and CsO molecules in Fig. 7 where most of the ECPs are well within the chemical accuracy regime. The ccECP shows noticeable improvements at the equilibrium bond length (Fig. 7). Similarly to the case of Rb, LANL2 has been omitted from the molecular binding curves due to its large discrepancies. Furthermore, for the ccECP we observe a systematic improvement in the atomic spectrum compared to the legacy ECPs, as shown in Figs. 1-3
@@
Although Ba has no occupied electrons in its ground state, the channel is crucial to capture subtle electronic interactions and relativistic effects, ensuring a more precise representation of this atom in complex systems. During optimization, we sought a balance between accurately replicating the AE atomic spectrum and maintaining robust performance across various chemical environments. Our primary focus was to prioritize precision in the atomic spectrum, accepting a minor trade-off in molecular binding curves (Fig. 8(b)) for BaO at compressed bond lengths. The ccECP demonstrates consistent results within chemical accuracy across a wide array of bond lengths, particularly at the equilibrium geometry. During the optimization step we focused on mitigating the overbinding issue at compressed bond lengths, achieving higher accuracy compared to several legacy ECPs and notably outperforming even the UC method in this critical region. In our view, this approach represents the optimal balance, maximizing the overall efficacy and reliability of the ccECP.
@@ III.3 Selected and elements
For antimony (Sb), there are nominally five valence electrons, making it rather difficult to capture accurate properties for all the tested possibilities. The results presented in Figs. 1-3 indicate that our ccECP performs on par or better than most legacy ECPs and the UC method. While it is possible to further improve the MADs through optimization, it comes at the cost of negatively impacting the accuracy of the binding curve in SbO. The optimal compromise is encouraging, compared to the legacy constructions; see Fig. 11, while also providing lower MADs and LMADs. This required significant interventions in the optimization process of the ccECP parameters, in particular, special attention was given to the and local channels where we systematically refined the coefficients of these channels by explicitly varying the first and second digits after the decimal point. The optimizer was configured to constrain the specific exponents of the local channel, keeping them unchanged while allowing the remaining parameters to be optimized. These hand-tuned adjustments were carefully tracked to assess their impact on the SbO binding curve to determine which parameter eliminates overbinding as seen in the other comparing ECPs. This resulted in discrepancies that were well below the chemical accuracy near equilibrium bond lengths 11.
@@
For lead (Pb), converging calculations at the CCSD(T) level required the use of unrestricted CCSD(T) and adjustments of the energy denominators associated with single and double excitations and the application of the direct inversion of iterative subspace technique in order to aid the convergence. In the atomic spectrum, the ccECP shows a significant improvement over legacy ECPs and the UC method, achieving the best overall metrics and approaching chemical accuracy.Molecular calculations (Fig. 12) proved to be more challenging for ECPs, particularly for PbO, where legacy ECPs and the UC method exhibit discrepancies almost an order of magnitude or more beyond the chemical accuracy. In PbH, our ccECP balances performance across bond lengths, addressing overbinding at the dissociation limit (seen in MDFSTU, MHFSTU, MWBSTU, and SBKJC) and underbinding at compressed bond lengths (observed with UC and CRENBL). For PbO, the ccECP demonstrates consistent improvement, maintaining chemical accuracy across most geometries and at the equilibrium configuration, while significantly reducing errors at compressed bond lengths.